OpenStudy (anonymous):

Jake forgot to replace the cap on a bottle of cologne. The cologne began to evaporate at a rate of 18% per day. If the original amount of the cologne in the bottle was 50 grams, I need to draw this graph that best represents the amount of cologne f(x) remaining in the bottle after x days

OpenStudy (anonymous):

Hm. Do you know what the equation would look like for this type of problem?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It's going to be\[y=50(0.18)^x\]where 50 is your starting amount, 0.18 is the percent it decreases and x is the days after Jake left the cap off. Do you know how an exponential graph looks?

OpenStudy (gorv):

50 is initial

OpenStudy (gorv):

first day 50*18/100=9 gram remaining=50-9=41 2nd day 41*18/100=7.38 ....... . . . .. so after x days remain = 50-50*(18)^x

OpenStudy (anonymous):

I am here, sorry

OpenStudy (anonymous):

So it will start at 50 and the next point is 41 drops to 7.38 so